# Persona-Augmented Benchmarking: Evaluating LLMs Across Diverse Writing Styles

### v1

@@ 3.3 Check Entailment of Rephrased Examples
We then verify that the modified examples retain all necessary information for accurate question answering. Initially, we manually reviewed 60 of the most extreme examples (i.e., those with the most substantial performance change from the original benchmark performance) for each benchmark. We then employ an LLM directly in our pipeline to determine whether the rephrased context entails all the ground-truth answers to the associated questions. To minimize errors, we evaluate the entailment capabilities of various LLMs and select those with the lowest false positive rates. We prioritize minimizing false positives over false negatives, as this conservative approach excludes examples lacking essential information, even if it means unnecessarily discarding some valid rephrased prompts. As a result, any performance degradation we observe represents a lower bound on the true impact of writing style variation, since all retained prompts are confirmed to preserve the necessary information.
@@ 3.4 Evaluate LLM Performance
We systematically evaluate LLM performance across three versions of each benchmark: the original examples, SAE-rephrased examples, and persona-based rephrased examples. We isolate the effects of the rephrasing process from those of introducing diverse writing styles by considering the former two versions as baselines.
@@ 4 Experiments
[…]
@@
LLMs
We use two open-weight LLMs, Gemma-3-27b-it and Qwen-3-32b, for rephrasing and entailment checking. This procedure produces two sets of rephrased prompts and ensures that any observed variation is not due to model selection. Both models show strong alignment when determining prompt entailment, with approximately 85% overlap across all three benchmarks. Therefore, we consolidate results and only show figures for Gemma-3-27b-it since it was more conservative when determining entailment. On the rephrased prompts, we evaluate the performance of ten LLMs for each set of experiments, covering four model families with varying model sizes and release dates: Gemma-3 (1b-it, 4b-it, 27b-it) (Team et al., 2025), Llama-3 (8b-instruct, 70b-instruct) (Grattafiori et al., 2024), Llama-4-scout-17b-16e-instruct with 109b parameters (Meta AI, 2025), Phi-4 (mini-instruct, -instruct, with 4B and 14B parameters respectively) (Abdin et al., 2024), Qwen-3 (8b, 32b) (Team, 2025). In subsequent sections, we refer to these models by their family name and size.
@@ 5 Findings
Though prior works found that LLMs tend to produce homogeneous text in open-ended generation tasks (e.g., essays, long-form QA, code generation) Alvero et al. (2024); Reinhart et al. (2025); Shypula et al. (2025), we find that LLM-rephrased prompts exhibit substantial linguistic variation when rephrasing with strict guidelines, i.e., instructed to maintain semantic content. To assess textual diversity we measure average distinct n-grams (across $n=2$ to $n=5$) and within-dataset cosine similarity between prompts. For a fair comparison with the original benchmark, we create a balanced subset of our augmented benchmark by randomly selecting one persona-based rephrasing per original prompt to match the original benchmark size. We repeat this process 25 times. We find that our persona-augmented subset exhibits higher diversity compared to the original benchmark with:
  • Higher distinct n-gram scores (0.840.840.84 compared to 0.750.750.75 with standard error <0.01<0.01< 0.01) suggesting greater lexical variety and reduced repetition of common phrases across rephrased prompts.
  • Lower average within-dataset cosine similarity between test instances (0 compared to 0.110.110.11 with standard error <0.01<0.01< 0.01), indicating that persona-based rephrasing increases the linguistic variation between independent prompts.
@@
While these aggregate ranges highlight overall model sensitivity, analyzing specific persona subsets provides more granular insights into writing style variations. Comparing best- and worst-performing persona subsets reveals average performance differences of 20, 11, and 28 percentage points for CoQA, CosmosQA, and DS-1000 respectively (Figure˜7). Qwen-3-32b exemplifies these effects with reductions from 0.14 to 0.12 (cosine similarity), 0.78 to 0.73 (accuracy), and 0.18 to 0.12 (accuracy) across benchmarks (Figure˜6). Some persona subsets trigger performance decrements up to 35% compared to SAE rephrasing.
[…]
We further examine how increasing writing style diversity affects the benchmark rankings. We hypothesize that writing style differences could potentially lead to instability in benchmark rankings in competitive leaderboards. Despite relatively stable rankings in our experiments (with only minor shifts, see Figure˜8)–likely stemming from our deliberate selection of models with widely varying capabilities–the magnitude of observed performance shifts (5-28 percentage points) could significantly impact rankings in scenarios with more closely matched models. Our simulations using the DS-1000 leaderboard4 indicate potential rank changes from -19 to +14 positions depending on the distribution of persona types used in evaluation (Figure˜9). [^5] This simulation assumes other models’ rankings would remain unchanged relative to their performance shifts. In such competitive environments, even a five percentage point performance shift could alter a model’s ranking by up to 16 positions. Similar patterns emerged in our CoQA and CosmosQA simulations.
@@ 6 Conclusion
Our research leverages sociodemographic attributes within personas to evaluate LLM performance across diverse writing styles. However, we emphasize that these synthetic personas cannot claim to authentically represent how specific demographic groups write in the real world. The personas serve as controlled variations to test model robustness, not as definitive representations of any particular population.
We recognize the ethical complexity of studying performance disparities across writing styles associated with different sociodemographic attributes. Our approach can be used to detect implicit biases that arise from an LLM’s perception of how certain groups communicate, without reinforcing harmful stereotypes. Specifically, we are examining how models respond to their own encoded assumptions about language variation and make no claims about actual communication patterns of different communities.
We deliberately avoid attributing poor model performance to specific demographic groups, focusing instead on identifying the linguistic patterns and writing styles that trigger performance disparities. This approach acknowledges the fluid nature of language and reduces the risk of stereotyping or stigmatizing particular populations.
We also recognize that our findings have implications for fairness in AI deployment. When certain writing styles consistently trigger lower model performance, this creates unequal access to AI capabilities across different user groups. By identifying these disparities, our research can be used to support more equitable LLM development that serves diverse linguistic communities effectively.
@@ Appendix F Analyzing Writing Styles
Table 6: Linguistic features for CoQA rephrased prompts compared the SAE rephrasing and the original benchmark. Thanks to the large sample size, the standard error of these estimates is <0.01<0.01< 0.01 across most metrics, making most of the differences statistically significant under Sign tests.
@@ Appendix G Variations Resulting from Base Personas
Appendix H Experimental Results
@@
Our experimental results reveal consistent and significant performance disparities between different persona subsets across all evaluated Large Language Models (LLMs). Figure˜7 quantifies these disparities, demonstrating substantial performance degradations of up to -23% for CoQA, -11% for CosmosQA, and -38% for DS-1000 when comparing the worst-performing persona subset (25th percentile) to the best-performing subset (75th percentile). Figure˜6 illustrates the absolute performance differences relative to the Standard American English (SAE) baseline, revealing that even the most recently released and advanced models like Llama-4-scout and Qwen-3-32b remain susceptible to writing style variations induced by different personas.
@@
As benchmarks are commonly used to select the best model for a task or compare model performance (such as on official benchmark leaderboards), we examine how ranking stability changes as we progress through different augmentation stages. Figure˜8 tracks ranking changes from the original benchmark through SAE rephrasing to increasingly challenging persona subsets (Best, All, Worst), with stability by Spearman correlation ($r$) and the Mann-Kendall rank correlation test ($\tau$) (Kendall, 1938).
@@
This stability breaks down in more competitive scenarios such as public or official benchmark leaderboards where models typically have much smaller performance gaps. As discussed in Section˜5, we simulated the DS-1000 leaderboard to examine how dense score distributions–typical of real leaderboards–affect ranking stability. Figure˜9 demonstrates that all models are highly sensitive to leaderboard changes depending on writing style. When comparing performance across different persona subsets (worst 25%, all, to best 25% of personas), models can shift by as much as -19 to +14 positions relative to their baseline rankings. These dramatic ranking changes—with some models experiencing swings of over 30 positions—stem from the substantial performance variations observed when models encounter different persona subsets. Such instability undermines the validity of current benchmarking practices and suggests that many performance differences may reflect sensitivity to writing style rather than true capability differences.
H.1 Correlation Between Model Performance
To understand how consistently different models respond to writing style variations, we analyze correlations between model performances across all 1200 personas. If models respond similarly to the same personas–both struggling with certain writing styles and excelling with others–we would expect high correlations. If models respond differently and unpredictably to the same personas, correlations would be low. Figures 10, 11, and 12 present correlation matrices showing both Pearson and Spearman rank correlations between all model pairs for each benchmark.
@@ H.1 Correlation Between Model Performance
Figure 10: Correlation of the performance on CoQA between all models on 1200 personas with average Pearson correlation coefficient $r=0.84$
Figure 11: Correlation of the performance on CosmosQA between all models on 1200 personas with average Pearson correlation coefficient $r=0.07$
Figure 12: Correlation of the performance on DS-1000 between all models on 1200 personas with average Pearson correlation coefficient $r=0.44$
H.2 Examining Persona Definitions Receiving the Worst Performance
@@
To identify which types of personas consistently cause performance degradation across models, we analyze performance patterns by sociodemographic attributes. Figure˜13 provides a detailed breakdown of performance changes by sociodemographic attributes, revealing particularly pronounced negative effects for specific attributes across all tasks and models.

### v2

@@ 3.3 Check Entailment of Rephrased Examples
We then verify that the modified examples retain all necessary information for accurate question answering. Initially, we manually reviewed 60 of the most extreme examples (i.e., those with the most substantial performance change from the original benchmark performance) for each benchmark. We then employ an LLM directly in our pipeline to determine whether the rephrased context entails all the ground-truth answers to the associated questions. To minimize errors, we evaluate the entailment capabilities of various LLMs and select those with the lowest false positive rates. We prioritize minimizing false positives over false negatives, as this conservative approach excludes examples lacking essential information, even if it means unnecessarily discarding some valid rephrased prompts. As a result, any performance degradation we observe represents a lower bound on the true impact of writing style variation, since all retained prompts are confirmed to preserve the necessary information. We confirm this intuition by employing a second entailment checking LLM. Using these two models, we obtain two ratings for entailment. If the models disagree on a particular example (i.e., one claims the answer is entailed by the context and the other does not), then we add this to the “disagreement region.” We use samples within this region to construct error bars for all our performance estimates to account for any biases exhibited by the entailment-checking LLM (see the full procedure in Appendix˜H).
@@ 3.4 Evaluate LLM Performance
We systematically evaluate LLM performance across three versions of each benchmark: original examples, SAE-rephrased examples, and persona-based rephrased examples. To isolate the effects of introducing diverse writing styles from a model having poor rephrasing abilities, we use both the original and SAE-rephrased examples as baselines. This design allows us to distinguish between performance degradation due to a poor rephrasing model (by comparing original to SAE-rephrased examples) and the effects of persona-induced writing styles (by comparing persona-based rephrased examples with SAE-rephrased examples).
@@ 4 Experiments
Figure 3: The performance changes (%) when using persona-rephrased prompts compared to the original benchmark across ten LLMs on three tasks: CoQA (top), CosmosQA (middle), and DS-1000 (bottom). Each violin plot represents the average performance difference received compared to SAE rephrasing by some LLM. The performance of all models are sensitive to writing styles with performance changes varying by 15-80%  [^3] Gemma-3-1b on DS-1000 is likely an outlier due to very low performance of $\approx 3\%$ accuracy even on the original benchmark. for a single model across different persona subsets.
@@
Entailment LLMs
We use two open-weight LLMs, Gemma-3-27b-it and Qwen-3-32b, for rephrasing and entailment checking. This procedure produces two sets of rephrased prompts and ensures that any observed variation is not due to model selection. Both models show strong alignment with each other when determining prompt entailment, with approximately 85% overlap across all three benchmarks. We consolidate results and show figures for Gemma-3-27b-it since it was more conservative in accepting entailment. To assess the sensitivity of the entailment model in our experimental setup, we use the 15% of cases where the models disagree on entailment to form our disagreement region, which is then used to construct error bars around our performance estimates. We find that across different subsets of personas, the worst performing personas (i.e., personas receiving performance in the 20th percentile) remain the same. Including any fraction of the disagreement region leads to worse results approximately 88% of the time. Thus any performance degradation reported in our paper is a conservative estimate. Additional details on these sensitivity analyses are provided in Appendix˜H.
Evaluation LLMs
On the rephrased prompts, we evaluate the performance of ten LLMs for each set of experiments, covering four model families with varying model sizes and release dates: Gemma-3 (1b-it, 4b-it, 27b-it) (Team et al., 2025), Llama-3 (8b-instruct, 70b-instruct) (Grattafiori et al., 2024), Llama-4-scout-17b-16e-instruct with 109b parameters (Meta AI, 2025), Phi-4 (mini-instruct, -instruct, with 4B and 14B parameters respectively) (Abdin et al., 2024), Qwen-3 (8b, 32b) (Team, 2025). In subsequent sections, we refer to these models by their family name and size.
@@ 5 Findings
Though prior works found that LLMs tend to produce homogeneous text in open-ended generation tasks (e.g., essays, long-form QA, code generation) Alvero et al. (2024); Reinhart et al. (2025); Shypula et al. (2025), we find that LLM-rephrased prompts exhibit substantial linguistic variation when rephrasing with strict guidelines, i.e., instructed to maintain semantic content.
To assess textual diversity we measure the average distinct n-grams (across $n=2$ to $n=5$) and within-dataset cosine similarity between prompts. For a fair comparison with the original benchmark, we create a balanced subset of our augmented benchmark by randomly selecting one persona-based rephrasing per original prompt to match the original benchmark size. We repeat this process 25 times. We find that our persona-augmented subset exhibits greater linguistic variation and reduced repetition of common phrases compared to the original benchmark as evidenced by higher distinct n-gram scores ($0.84$ compared to $0.75$ with standard error <0.01<0.01). Additionally, persona-based rephrasing increases the linguistic variation between independent prompts with lower average within-dataset cosine similarity between test instances (0 compared to $0.11$ with standard error <0.01<0.01).
@@
While these aggregate ranges highlight overall model sensitivity, analyzing specific persona subsets provides more granular insights into writing style variations. Comparing best- and worst-performing persona subsets reveals average performance differences of 20, 11, and 28 percentage points for CoQA, CosmosQA, and DS-1000 respectively (Figure˜8). Qwen-3-32b exemplifies these effects with reductions from 0.14 to 0.12 (cosine similarity), 0.78 to 0.73 (accuracy), and 0.18 to 0.12 (accuracy) across benchmarks (Figure˜6). Some persona subsets trigger performance decrements up to 35% compared to SAE rephrasing.
Sensitivity to linguistic variations appears to be correlated with task complexity. The most challenging benchmark, DS-1000 (with $59\%$ accuracy from state-of-the-art models [^4] The official leaderboard can be found at https://ds1000-code-gen.github.io), exhibits the widest performance spread, while the least challenging one, CosmosQA, shows little variation. This may suggest that commonsense reasoning is less sensitive to linguistic variation compared to conversational question-answering and code generation tasks.
We further examine how increasing writing style diversity affects the benchmark rankings. We hypothesize that writing style differences could potentially lead to instability in benchmark rankings in competitive leaderboards. Despite relatively stable rankings in our experiments (with only minor shifts, see Figure˜9)–likely stemming from our deliberate selection of models with widely varying capabilities–the magnitude of observed performance shifts (5-28 percentage points) could significantly impact rankings in scenarios with more closely matched models. Our simulations using the DS-1000 leaderboard4 indicate potential rank changes from -19 to +14 positions depending on the distribution of persona types used in evaluation (Figure˜10). [^5] This simulation assumes other models’ rankings would remain unchanged relative to their performance shifts. In such competitive environments, even a five percentage point performance shift could alter a model’s ranking by up to 16 positions. Similar patterns emerged in our CoQA and CosmosQA simulations.
@@ 6 Conclusion
While our research leverages personas with varying sociodemographic attributes to evaluate LLM performance, we emphasize that these personas are not meant to represent how specific demographic groups write in the real world. The personas serve as a tool to create more diverse data to test model robustness and not as definitive representations of any particular population. Therefore, our analysis focuses on how models respond to their own encoded assumptions about language variation.
This approach can then be used to detect implicit biases resulting from a model’s perception without perpetuating harmful stereotypes about particular populations. This perception is shaped by the model’s training data and fine-tuning procedures. We instead report specific linguistic features and stylistic elements that trigger performance disparities rather than attributing variations to specific demographic groups. This approach acknowledges the dynamic and contextual nature of language use while reducing the risk of stigmatizing or stereotyping particular communities.
Our research findings also have implications for fairness in AI deployment. When certain writing styles consistently produce lower model performance, this creates differential access to AI capabilities across user populations. Researchers and practitioners can use our methodology to support more equitable LLM development that serves diverse linguistic communities effectively.
@@ Appendix F Analyzing Writing Styles
Table 6: Linguistic features for CoQA rephrased prompts compared the SAE rephrasing and the original benchmark. Thanks to the large sample size, the standard error of these estimates is <0.01<0.01 across most metrics, making most of the differences statistically significant under Sign tests.
@@ Appendix G Variations Resulting from Base Personas
Appendix H Accounting for Errors in Entailment-Checking
We investigate and quantify the sensitivity of persona-augmented benchmarking to entailment model bias or error. We estimate performance for evaluation sets containing only prompts where both models agree on entailment, then systematically vary the proportion of disputed prompts to assess sensitivity to entailment model disagreements. We then report error bars for performance changes rather than strict point estimates.
H.1 Methodology
We assess whether the results reported in the main body of our paper would change when our test dataset includes cases where two entailment models disagree. We define the agreement region as prompts where both Gemma-3-27b-it and Qwen-3-32b-instruct reach the same entailment decision, and the disagreement region as prompts where they disagree on entailment. We systematically vary the proportion of disagreement region data (25%, 50%, 75%, 100%) included alongside the agreement region data to create error ranges for each persona’s performance estimate. This process is repeated 10 times with reshuffled data from the disagreement region. This procedure identifies the maximum uncertainty range that disagreement cases could introduce. This method greedily returns minimum and maximum bounds indicating the range of performance degradation when varying evaluation samples. We can use these bounds to assess whether disagreements between entailment models systematically bias our performance estimates.
Figure 7: Percent change in performance across all three datasets with error bars on each model representing the range of possible performance estimates when entailment models disagree.
H.2 Results
We find that entailment model uncertainty primarily results in worse performance of about 4 percentage points, while on average improving performance by about 2 percentage points. Even with this variation, our findings show significant performance degradation for different persona subsets (as shown by the best compared to worst personas for CoQA and CosmosQA in Figure 7). The systematic performance differences between these persona subsets remain statistically significant across all evaluation configurations. Additionally, since point estimates consistently fall near the top of the error bars rather than the center, these results support our hypothesis that using conservative entailment filtering creates a lower bound on true performance degradation.
Our core findings are robust to this variation, showing substantial performance drops persist across different persona subsets, and the worst-performing personas remain consistent regardless of entailment model choice or evaluation set composition. This consistency demonstrates that regardless of which entailment model is used and how conservative our approach is, our overall findings remain the same. There is still a substantial drop in performance across different subsets of personas, and the worst-performing personas often remain the same.
Appendix I Experimental Results
@@
Our experimental results reveal consistent and significant performance disparities between different persona subsets across all evaluated Large Language Models (LLMs). Figure˜8 quantifies these disparities, demonstrating substantial performance degradations of up to -23% for CoQA, -11% for CosmosQA, and -38% for DS-1000 when comparing the worst-performing persona subset (25th percentile) to the best-performing subset (75th percentile). Figure˜6 illustrates the absolute performance differences relative to the Standard American English (SAE) baseline, revealing that even the most recently released and advanced models like Llama-4-scout and Qwen-3-32b remain susceptible to writing style variations induced by different personas.
@@
As benchmarks are commonly used to select the best model for a task or compare model performance (such as on official benchmark leaderboards), we examine how ranking stability changes as we progress through different augmentation stages. Figure˜9 tracks ranking changes from the original benchmark through SAE rephrasing to increasingly challenging persona subsets (Best, All, Worst), with stability by Spearman correlation ($r$) and the Mann-Kendall rank correlation test ($\tau$) (Kendall, 1938).
@@
This stability breaks down in more competitive scenarios such as public or official benchmark leaderboards where models typically have much smaller performance gaps. As discussed in Section˜5, we simulated the DS-1000 leaderboard to examine how dense score distributions–typical of real leaderboards–affect ranking stability. Figure˜10 demonstrates that all models are highly sensitive to leaderboard changes depending on writing style. When comparing performance across different persona subsets (worst 25%, all, to best 25% of personas), models can shift by as much as -19 to +14 positions relative to their baseline rankings. These dramatic ranking changes—with some models experiencing swings of over 30 positions—stem from the substantial performance variations observed when models encounter different persona subsets. Such instability undermines the validity of current benchmarking practices and suggests that many performance differences may reflect sensitivity to writing style rather than true capability differences.
I.1 Correlation Between Model Performance
To understand how consistently different models respond to writing style variations, we analyze correlations between model performances across all 1200 personas. If models respond similarly to the same personas–both struggling with certain writing styles and excelling with others–we would expect high correlations. If models respond differently and unpredictably to the same personas, correlations would be low. Figures 11, 12, and 13 present correlation matrices showing both Pearson and Spearman rank correlations between all model pairs for each benchmark.
@@ I.1 Correlation Between Model Performance
Figure 11: Pearson correlation between all models’ performance on CoQA over 1200 personas ($r=0.84$)
Figure 12: Pearson correlation between all models’ performance on CoQA over 1200 personas ($r=0.07$)
Figure 13: Pearson correlation between all models’ performance on CoQA over 1200 personas ($r=0.44$)
I.2 Examining Persona Definitions Receiving the Worst Performance
@@
To identify which types of personas consistently cause performance degradation across models, we analyze performance patterns by sociodemographic attributes. Figure˜14 provides a detailed breakdown of performance changes by sociodemographic attributes, revealing particularly pronounced negative effects for specific attributes across all tasks and models.